OpenStudy (anonymous):

Solve 5x^3-6x^2-4x-8=0

OpenStudy (anonymous):

plug x=2 f(x)=5(2)^3-6x^2-4x-8 =40-24 -8-8=0 hence x=2 is a root of f(x)=0 factorise by synthetic division with 2

OpenStudy (anonymous):

lol don't tell me what to do, I do what I want. I think surjithayer deserves the medal for not demanding it. lol